Why do we measure perimeter in ft, area in ft^2, and volume in ft^3?
What does each mean within the strand of measurement?
Why do students label with the incorrect exponent?

Answer by mananth(16949)

Why do we measure perimeter in ft, area in ft^2, and volume in ft^3?
What does each mean within the strand of measurement?
Why do students label with the incorrect exponent?

I will try to explain with an example
The length is a linear measurement.
it is measured in inches,feet or yards or miles.
For short lengths we generally measure in inches and feet,
The bench is 3 feet in length
The pencil is eight inches in length.
The distance between two cities is ,measured in miles. Because the measurement of large distances in feet or inches involves very big numbers.
Length is measured in single units.

Area ;
Area involves the flat surface enclosed by figure having three or more side.

for example area of a rectangle is length in feet * width in feet
8feet * 6 feet . Actually you are multiplying the digits 8and 6 so also you are multiplying feet.
ie: 8feet * 6 feet = 8*6* feet*feet = 48 feet square or commonly square feet.

when volume is involved there are three dimensions involved.
Volume of a cuboid is given by length * width* height
example = cuboid of dimensions 6feet by 7 feet by 8 feet
we write 6feet * 7 feet*8 feet.
it is actually 6*7*8* feet*feet*feet = 336 feet cubic or commonly cubic feet.
Whatever may be the units , length is always in single unit, area in square units and vloume in cubic units